# FRWKV: FREQUENCY-DOMAIN LINEAR ATTENTION FOR LONG-TERM TIME SERIES FORECASTING

### v1

@@ 2.1 Problem Definition
Given a multivariate time series $\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}]\in\mathbb{R}^{N\times T}$, where NN is the number of variables and $T$ is the sequence length, the long-term forecasting task aims to predict future values $\mathbf{Y}=[\mathbf{x}_{T+1},\mathbf{x}_{T+2},\ldots,\mathbf{x}_{T+\tau}]\in\mathbb{R}^{N\times\tau}$ for a horizon $\tau$. Our goal is to learn a mapping function $f:\mathbb{R}^{N\times T}\to\mathbb{R}^{N\times\tau}$ that minimizes the prediction error while maintaining computational efficiency.
@@ 2.2 Model Overview
We use the following symbols in the paper and Fig. 1: $B$ denotes batch size; $T$ is the input sequence length (time steps); NN is the number of variables; $E$ is the embedding dimension; $C$ is the model dimension (d_model); $F=\lfloor T/2\rfloor+1$ is the number of frequency bins after rFFT; $H$ is the number of attention heads and $D_{h}$ is the attention head size, satisfying $H\cdot D_{h}=C$.

### v2

@@ 2.1 Problem Definition
Given a multivariate time series $\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}]\in\mathbb{R}^{N\times T}$, where $N$ is the number of variables and $T$ is the sequence length, the long-term forecasting task aims to predict future values $\mathbf{Y}=[\mathbf{x}_{T+1},\mathbf{x}_{T+2},\ldots,\mathbf{x}_{T+\tau}]\in\mathbb{R}^{N\times\tau}$ for a horizon $\tau$. Our goal is to learn a mapping function $f:\mathbb{R}^{N\times T}\to\mathbb{R}^{N\times\tau}$ that minimizes the prediction error while maintaining computational efficiency.
@@ 2.2 Model Overview
We use the following symbols in the paper and Fig. 1: $B$ denotes batch size; $T$ is the input sequence length (time steps); $N$ is the number of variables; $E$ is the embedding dimension; $C$ is the model dimension (d_model); $F=\lfloor T/2\rfloor+1$ is the number of frequency bins after rFFT; $H$ is the number of attention heads and $D_{h}$ is the attention head size, satisfying $H\cdot D_{h}=C$.